# TOTEM: TOkenized Time Series EMbeddings for General Time Series Analysis

### v1

@@ 3 Method
Our proposed discrete time series tokenization enables the design of general models across a variety of time series domains, tasks, and evaluation schemas, Figure 1. We design a single tokenizer architecture that is generally applicable without extensive data engineering while being suitable for varying data dimensionalities across different tasks. There are many possibilities for how to introduce a discrete time series tokenizer, we extensively study one such methodology that satisfies the aforementioned design criteria.
Data Engineering. Prior work in time series analysis leverages data engineering such as the use of auxiliary features (e.g. day of the month, or minute in the hour, etc.) (Chen et al., 2023; Salinas et al., 2020), or frequency transformations (Wu et al., 2022a; Zhou et al., 2022). We forego any data engineering and operate directly on time steps. This enables generalist-training as differing data domains have widely varying sampling rates leading to distinct auxiliary features and frequency profiles.
Varying Dimensionality. A time series dataset consists of $E$ examples (i.e. number of distinct recordings), $S$ sensor channels, and $T$ time steps, and can be formally expressed as $\{\mathbf{x}_{j}\}_{j=1}^{E}\subset\mathbb{R}^{S\times T}$. Even within a single task and single data domain where $S$ does not change, $E$ and $T$ take on a wide range of values. As an example, canonical forecasting predictions lengths range from $96$ to $720$ time steps. When moving to generalist-training, datasets additionally have wide ranging sensor dimensionalities $S$. Our tokenizer handles varying dimensionality across $E$, $S$, and $T$ by creating non-overlapping tokens along the time-dimension that are smaller than the dimension $T$.
Differing Tasks. There are numerous tasks to tackle in time series analysis. Three significant ones are imputation, anomaly detection, and forecasting. In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$, and then reconstruct and impute $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$. In anomaly detection, models intake a corrupted time series $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and reconstruct the data $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$. The amount of corruption is considered known, at A%. In forecasting, models intake a time series $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict future readings $\mathbf{y}\in\mathbb{R}^{S\times T_{\textrm{out}}}$, where $S$ is the number of sensors and $T_{\textrm{in}},T_{\textrm{out}}$ signify the durations of the preceding and succeeding time series, respectively. Our tokenizer is performant across all tasks despite their distinct representational requirements.
TOTEM Implementation. To realize a single tokenizer architecture that enables generalist modeling across differing domains and tasks we take inspiration from the VQVAE (Van Den Oord et al., 2017). The original VQVAE leverages a dilated convolutional architecture with a stride of 2 and window-size of 4, similar to the WaveNet (Oord et al., 2016) dilated, causal, convolutional decoder. A dilated convolution skips inputs allowing a filter to operate on a larger input area / coarser scale. Utilizing dilated convolutions is an architectural decision rooted in the high sampling rates of raw audio waveforms (Oord et al., 2016; Van Den Oord et al., 2017). High sampling rates are not a trait shared by many time series domains.
When adapting the VQVAE for general time series analysis, the TOTEM VQVAE:
  1. 1.
    Operates directly on time steps; no data engineering.
  2. 2.
    Creates discrete, non-overlapping tokens along the time dimension of length $F$, where $F<T$, thereby promoting training and testing on variable length examples, $E$, sensors, $S$, and time steps $T$.
  3. 3.
    Maintains the same architecture and objective regardless of the downstream task.
  4. 4.
    Aims to capture maximal information within a large receptive field by: (1) using a strided non-causal convolutional architecture with no dilation, (2) training on long time series inputs, (3) pre-striding the data by a stride of 1 so the tokenizer learns from maximal inputs.
The TOTEM VQVAE consists of an encoder, quantizer, latent codebook, and decoder. It takes in a univariate time series $\{\mathbf{x}_{i}\in\mathbb{R}^{T}\}_{i=1}^{E\cdot S}$ obtained by flattening the sensor channel of the multivariate data. This makes TOTEM’s VQVAE sensor-agnostic, enabling TOTEM’s generalist-training and zero-shot-testing. The encoder $\mathcal{E}$ consists of strided 1D convolutions compressing the time series by a cumulative stride of $F$. $\mathcal{E}$ maps a univariate time series $\mathbf{x}\in\mathbb{R}^{T}$ to a latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{\nicefrac{{T}}{{F}}\times D}$, where $D$ is the the hidden dimension. The latent codebook $\mathcal{C}=\{\mathbf{c}_{i}\}_{i=1}^{K}$ consists of $K$ $D$-dim codewords $\mathbf{c}_{i}\in\mathbb{R}^{D}$. During quantization, the codebook is used to replace $\mathbf{z}$ with $\hat{\mathbf{z}}\in\mathbb{R}^{\nicefrac{{T}}{{F}}\times D}$ such that $\hat{\mathbf{z}}_{j}=\mathbf{c}_{k}$, where $k=\operatorname*{arg\,min}_{i}||\mathbf{z}_{j}-c_{i}||_{2}$. The decoder $\mathcal{D}$ follows the reverse architecture of the encoder $\mathcal{E}$, consisting of 1D transpose convolutions with a cumulative stride of $\nicefrac{{1}}{{F}}$ mapping the quantized $\hat{\mathbf{z}}$ to a reconstructed time series $\hat{\mathbf{x}}=\mathcal{D}(\hat{\mathbf{z}})\in\mathbb{R}^{T}$. We learn $\mathcal{E}$, $\mathcal{D}$, and $\mathcal{C}$ by optimizing the objective $\mathcal{L}=\mathcal{L}_{\textrm{rec}}+\mathcal{L}_{\textrm{cmt}}$ consisting of a reconstruction loss $\mathcal{L}_{\textrm{rec}}~{}=~{}\frac{1}{E\cdot S}\sum_{i}||\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}||^{2}_{2}$ and a commitment loss $\mathcal{L}_{\textrm{cmt}}$, which allows the codebook to update despite the the non-differentiable $\operatorname*{arg\,min}$ operation during quantization. The final objective is $\mathcal{L}=\mathcal{L}_{\textrm{rec}}+\alpha\cdot\mathcal{L}_{\textrm{cmt}}$, where $\alpha$ is a scalar that weights the two losses. This objective does not change even when the underlying task, time series length, data masking, normalization schema, or data domain changes.
Notably imputation and anomaly detection can be directly solved with just TOTEM’s VQVAE, as they are fundamentally data representation tasks, whereas in forecasting further modeling is required, see Figure 2. In forecasting, the trained, frozen, codebook representation converts a sensor’s observed measurements $\mathbf{x}_{s}\in\mathbb{R}^{T_{\text{in}}}$ to a sequence of $\nicefrac{{T_{\text{in}}}}{{F}}$ discrete tokens. The forecaster transformer encoder processes these tokenized time series independently for each sensor, adding time-based positional encodings to each token along the time dimension. Using a series of multi-head attention layers, the model predicts the forecasted measurements $\bar{\mathbf{y}}_{s}\in\mathbb{R}^{T_{\textrm{out}}}$ for $s=1,...,S$, applying the attention mechanism along the time dimension $T$. In parallel, the forecaster takes in $\mathbf{x}_{s}$ and predicts the future’s mean, $\mu_{s}$, and standard deviation, $\sigma_{s}$, for each sensor $s=1,...,S$ to unnormalize the data. The final forecasted prediction is $\mathbf{y}_{s}=\sigma_{s}\cdot\bar{\mathbf{y}}_{s}+\mu_{s}$. The forecaster is trained in a supervised fashion by minimizing three smooth L1 losses between predictions $\{\bar{\mathbf{y}}_{s},\mu_{s},\sigma_{s}\}$ and their ground truth respectively.
Table 1: Specialist Imputation ()normal-↓($\downarrow$)( ↓ ). Across all datasets, metrics, and masking percentages, TOTEM has the highest AvgWins (52.1%), followed by GPT2 (35.4%). TOTEM values are means from 3 seeds; baseline values are from (Zhou et al., 2023; Wu et al., 2022a).

### v2

@@ 3 Method
3.1 Task Definitions
This work considers three tasks: imputation, anomaly detection, and forecasting. In imputation, models intake a masked time series $\mathbf{x_{m}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and impute the missing values to recover the reconstruction $\mathbf{\hat{x}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$. In anomaly detection, models intake a time series corrupted at a known level $\mathbf{x_{corr}}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict which times are anomalous, $\mathbf{y}\in\{0,1\}^{T_{\text{in}}}$. Lastly, in forecasting, models intake a time series $\mathbf{x}\in\mathbb{R}^{S\times T_{\textrm{in}}}$ and predict future values $\mathbf{y}\in\mathbb{R}^{S\times T_{\textrm{out}}}$, where $T_{\textrm{in}}$ and $T_{\textrm{out}}$ signify the durations of the preceding and succeeding time series, respectively. A core design goal for TOTEM is to learn a representation suitable for any of these three tasks using the same architecture and without leveraging any task- or domain-specific knowledge.
3.2 Design Decisions
This section discusses TOTEM’s key design features: a self-supervised training stage, exclusively-temporal tokenization, and no domain-specific data engineering.
Self-supervised Tokenizer Training. As described in Section 2, TOTEM learns a fixed codebook of tokens over a multi-domain corpus of time series data independently from the training of any downstream model. This disentangles the choice of data representation from the choice of task-specific architecture and permits the learning of representations from a large, diverse set of data, which aids in zero-shot generalization.
First, we elect to use a discrete, deterministic encoder to produce time series tokens. This decision is largely motivated by large language models (and in particular, tokenization methods in NLP like byte pair encoding (BPE) (Gage, 1994; Radford et al., 2018)), in which a downstream model learns on a finite number of distinct tokens. Moreover, in methods like BPE, the tokenization operation is lossless and reversible because it is deterministic (though non-unique). This suggests that vector quantization-based models could be effective for tokenizing time series data. Two popular vector quantization methods are VQVAEs (Van Den Oord et al., 2017) and VQGANs (Esser et al., 2021). In this work, we choose to use a VQVAE, as VQGANs are more commonly used for encoding images. Moreover, the use of VQVAEs has been studied in neural audio models (Oord et al., 2016; Van Den Oord et al., 2017), including followup works with audio-specific models (Baevski et al., 2020), which suggests that they may be effective for modeling general time series.
Figure 2: Left. Specialist models can tokenize along any of the $E$, $S$, or $T$ dimensions. Right. Generalist models can only tokenize along $T$, since the learned tokenization must apply to a diverse set of domains with any possible data dimensionality.
Exclusively-Temporal Tokenization. A time series dataset consists of $E$ examples, $S$ sensor channels, and $T$ time steps, and can be formally expressed as $\{\mathbf{x}_{j}\}_{j=1}^{E}\subset\mathbb{R}^{S\times T}$. Prior work commonly patches along either the sensor dimension (Li et al., 2019; Zhou et al., 2021; Wu et al., 2021; Liu et al., 2021), or time dimension (Liu et al., 2023; Zhang & Yan, 2022; Nie et al., 2022). When training specialists, it is reasonable to tokenize across any combination of these or the example dimension (e.g., in neuroscience data, it is common to group recordings by day, where the subject exhibits different behavior on a daily basis (Talukder et al., 2022)).
However, in the generalist case, because the sensors associated with each domain have distinct semantic meanings, performing sensor- or example-wise tokenization will capture domain-specific relations, hindering the tokenizer’s generalization. Thus, we choose to exclusively tokenize over the temporal dimension, such that the tokens represent univariate waveforms. Further, this is crucial for testing the tokenizer in the zero-shot regime, where the dimensionality of the testing domain may differ significantly from that of the training domain(s). Specifically, TOTEM tokenizes time series data with non-overlapping temporal chunks of length $T/F$, where $F$ is some compression factor for downsampling the data.
No Domain-specific Data Engineering. Many prior works (especially in time series forecasting) leverage domain-specific knowledge to handcraft features that encode critical information. For instance, works that study calendar-based time series often add auxiliary features that denote landmarks like the first day of a month or holidays (Chen et al., 2023; Salinas et al., 2020). Other works propose highly-engineered architectures that convert time series into frequency space representations. For example, TimesNet operates on the assumption that most time series exhibit multi-resolutional periodicity, and convert a time series into a frequency-space image by computing the Fourier transform on several subsets of the time series (Wu et al., 2022). Similarly, FedFormer represents a time series with a random subset of its Fourier components and a complex mixture-of-experts model (Zhou et al., 2022). In contrast, TOTEM uses only reverse instance normalization (RevIN) (Kim et al., 2021) to represent temporal waveforms in a normalized space (see Figure 3), which requires no assumptions on the form of the data. This allows TOTEM to generalize across domains and outperform the prior handcrafted methods on many distinct tasks using simple, generic architectures.
3.3 Tokenizer Implementation
Figure 3: TOTEM flattens the sensor and example dimensions and learns a discrete representation along the time dimension in a normalized space.
Though TOTEM is a VQVAE, the design of the encoder and decoder differ substantially from the original model and similar works like WaveNet, which use dilated convolutions (Oord et al., 2016; Van Den Oord et al., 2017). The dilations in these architectures skip many time steps, allowing the convolutional filters to operate on a larger input area at a coarser scale, improving model efficiency. However, this design decision is motivated by the high sampling rates of digital audio waveforms, which is not a universal trait across time series domains (see Table 6). In contrast, TOTEM uses a stack of strided 1D convolutions with a dilation of 1 such that it can account for every time step. Using a long input (e.g., 96 time steps for standard forecasting tasks) allows TOTEM to maintain a large receptive field. Lastly, the use of RevIN allows TOTEM to remain effective by only learning a small set of normalized waveforms, and if the unnormalized reconstruction is required for a downstream task, the normalization parameters can also be passed to the decoder (see Figure 4).
Formally, TOTEM accepts a batch of univariate time series $\{\mathbf{x}_{i}\in\mathbb{R}^{T}\}_{i=1}^{E\cdot S}$ obtained by flattening the sensor channel of the multivariate data. An encoder $\mathcal{E}$ consisting of a stack of strided 1D convolutions then temporally compresses the data by a total factor of $F$ to recover a latent variable $\mathbf{z}=\mathcal{E}(\mathbf{x})\in\mathbb{R}^{\nicefrac{{T}}{{F}}\times D}$, where $D$ is the latent feature dimension. The latent variable $\mathbf{z}$ is then quantized into an element $\mathbf{\hat{z}}$ of the codebook $\mathcal{C}=\{\mathbf{c}_{i}\}_{i=1}^{K}$ consisting of $K$ $D$-dimensional codewords $\mathbf{c}_{i}\in\mathbb{R}^{D}$ following the relation $\hat{\mathbf{z}}=\mathbf{c}_{\ell}$, where $\ell=\operatorname*{arg\,min}_{i}||\mathbf{z}-c_{i}||_{2}^{2}$. The decoder $\mathcal{D}$ mirrors the encoder’s architecture, mapping the quantized embedding $\hat{\mathbf{z}}$ to a reconstructed time series $\hat{\mathbf{x}}=\mathcal{D}(\hat{\mathbf{z}})\in\mathbb{R}^{T}$.
As in Van Den Oord et al. (2017), we train $\mathcal{E}$, $\mathcal{D}$, and $\mathcal{C}$ by optimizing the objective
$\mathcal{L}=\underbrace{\frac{1}{E\cdot S}\sum_{i}||\mathbf{x}_{i}-\hat{\mathbf{x}}_{i}||^{2}_{2}}_{\mathcal{L}_{\text{rec}}}+\underbrace{||\texttt{sg}[\mathbf{z}]-\mathbf{\hat{z}}||_{2}^{2}}_{\mathcal{L}_{\text{vq}}}+\beta\underbrace{||\mathbf{z}-\texttt{sg}[\mathbf{\hat{z}}]||_{2}^{2}}_{\mathcal{L}_{\text{cmt}}},$ (1)
where $\texttt{sg}[\cdot]$ is the stop-gradient operator and $\beta$ is the commitment loss weight. For additional details, see Appendices A.11 and A.12. In all experiments, we use a compression factor of $F=4$, (see Table 32).
3.4 Forecasting Model Implementation
In contrast with prior works, TOTEM is capable of solving the imputation and anomaly detection tasks with the tokenizer alone (see Figures 11 and 12). Therefore, the only downstream model we must design is the forecasting model. First, each sensor’s observations $\mathbf{x}_{i}\in\mathbb{R}^{T_{\text{in}}}$ are converted into a sequence of $\nicefrac{{T_{\text{in}}}}{{F}}$ discrete tokens $\mathbf{\hat{z}}_{i}$. The forecaster processes adds temporal positional embeddings to these tokens, passing them through a transformer encoder consisting of a series of multi-head attention layers that attend along the time dimension to predict normalized measurements $\bar{\mathbf{y}}_{i}\in\mathbb{R}^{T_{\textrm{out}}}$ for $i=1,...,S$.
Figure 4: The Forecaster Model. The forecaster takes in a tokenized version of normalized time series observations (obtained using TOTEM’s encoder) and predicts a normalized time series over some specified horizon along with parameters that allow the model to unnormalize the prediction.
A separate prediction head predicts the mean $\mu_{i}$ and standard deviation $\sigma_{i}$ associated with each univariate time series $\mathbf{x}_{i}$ such that the final forecasted prediction is $\mathbf{y}_{i}=\sigma_{i}\cdot\bar{\mathbf{y}}_{i}+\mu_{i}$ for $i=1,\dots,S$. The forecaster is trained in a supervised fashion by minimizing three smooth L1 losses between predictions $\{\bar{\mathbf{y}}_{i},\mu_{i},\sigma_{i}\}_{i=1}^{S}$ and their ground truth values respectively. Crucially, this architecture is used for all domains in our forecasting experiments, demonstrating that TOTEM can competitively perform forecasting in a domain-agnostic manner.
4 Experimental Setup
This section explains the experimental setup for each task, including the baselines and datasets used for evaluation. The results and analyses are presented in Section 5. We compare to two families of approaches: methods designed for multiple tasks (multi-task), like TOTEM, and methods designed for a specific task (single-task). Many single-task methods have frequently been adapted by others to tasks besides the ones for which they were originally designed, and in those cases, we compare against the best reported results for the adapted model. For all tasks, we trained a GPT2 generalist baseline from scratch, and for forecasting, we additionally trained a GPT2 specialist.
4.1 Imputation
Baselines. In the main text, we compare TOTEM against 12 baselines with varying model architectures. We further compare against 5 additional baselines with different architectures for completeness in the Appendix A.2. In total, we evaluate against 17 baselines. See Table 2A for a summary.
Datasets. For the in-domain testing regime, we test on 6 datasets, and for the zero-shot testing regime, we evaluate on an additional 5 datasets. We also perform additional evaluations in the appendix on the PhysioNet Challenge 2012 dataset. In total, we evaluate on 12 distinct datasets. See Table 2B for a summary.
Metrics. We report the mean squared error (MSE) and mean absolute error (MAE) of the imputed versus the ground truth signals.
A. Imputation Baselines
Type
Model
Abbr.
Arch.
Citation
Multi-task
GPT2 - Generalist
GPT2
TF
Trained by us
GPT2 - Specialist
GPT2
TF
Zhou et al. (2023)
TimesNet
TiNet
Conv
Wu et al. (2022)
Single-task
PatchTST
Patch
TF
Nie et al. (2022)
ETSFormer
ETS
TF
Woo et al. (2022)
Fedformer
FED
TF
Zhou et al. (2022)
Non-stationary Trans.
Stat
TF
Liu et al. (2022b)
Autoformer
Auto
TF
Wu et al. (2021)
Informer
Inf
TF
Zhou et al. (2021)
Reformer
Re
TF
Kitaev et al. (2020)
LightTS
LiTS
Linear
Zhang et al. (2022)
DLinear
DLin
Linear
Zeng et al. (2023)
Appendix
V-RIN
VAE
Mulyadi et al. (2021)
BRITS
RNN
Cao et al. (2018)
RDIS
Diffusion
Choi et al. (2023)
Unconditional CSDI
Diffusion
Tashiro et al. (2021)
CSDI
Diffusion
Tashiro et al. (2021)
B. Imputation Datasets
Regime
Dataset
Abbr.
Citation
In-Domain
Weather
W
Zhou et al. (2023)
Electricity
E
Zhou et al. (2023)
ETTm1
m1
Zhou et al. (2023)
ETTm2
m2
Zhou et al. (2023)
ETTh1
h1
Zhou et al. (2023)
ETTh2
h2
Zhou et al. (2023)
Zero-Shot
Neuro2
N2
Peterson et al. (2022)
Neuro5
N5
Peterson et al. (2022)
Saugeen River Flow
R
Godahewa et al. (2021)
U.S. Births
B
Godahewa et al. (2021)
Sunspot
S
Godahewa et al. (2021)
Appendix
PhysioNet
Silva et al. (2012)
Table 2: Imputation baselines and datasets.
4.2 Anomaly Detection
Baselines. In the main text, we compare TOTEM against 16 baselines, and in the Appendix A.3 an additional 3 for a total of 19 baselines (see Table 13). See Table 3A for a summary.
Datasets. For the in-domain testing regime, we test on 5 datasets, and for the zero-shot regime, we test on another 5. For additional signal, we also test on 15 distinct anomaly detection datasets from Wu & Keogh (2021) in Appendix A.3 (see Table 13). In total, we evaluate on 25 datasets. See Table 3B for a summary.
Metrics. We report the precision (P), recall (MSE), and adjusted F1 score.
A. Anomaly Detection Baselines
Type
Model
Abbr.
Arch.
Citation
Multi-task
GPT2 - Generalist
GPT2
TF
Trained by us
GPT2 - Specialist
GPT2
TF
Zhou et al. (2023)
TimesNet
TiNet
Conv
Wu et al. (2022)
Single-task
Anomaly Trans.
ATran
TF
Xu et al. (2021)
PatchTST
Patch
TF
Nie et al. (2022)
ETSFormer
ETS
TF
Woo et al. (2022)
Fedformer
FED
TF
Zhou et al. (2022)
Non-stationary Trans.
Stat
TF
Liu et al. (2022b)
Autoformer
Auto
TF
Wu et al. (2021)
Pyraformer
Pyra
TF
Liu et al. (2021)
Informer
Inf
TF
Zhou et al. (2021)
Reformer
Re
TF
Kitaev et al. (2020)
LogTrans
LogTr
TF
Li et al. (2019)
Transformer
Trans
TF
Vaswani et al. (2017)
LightTS
LiTS
Linear
Zhang et al. (2022)
DLinear
DLin
Linear
Zeng et al. (2023)
Appendix
DGHL
Challu et al. (2022)
MOMENT-0
Goswami et al. (2024)
MOMENT-LP
Goswami et al. (2024)
B. Anomaly Detection Datasets
Regime
Dataset
Abbr.
Citation
In-Domain
SMD
Zhou et al. (2023)
MSL
Zhou et al. (2023)
SMAP
Zhou et al. (2023)
SWAT
Zhou et al. (2023)
PSM
Zhou et al. (2023)
Zero-Shot
Neuro2
N2
Peterson et al. (2022)
Neuro5
N5
Peterson et al. (2022)
Saugeen River Flow
R
Godahewa et al. (2021)
U.S. Births
B
Godahewa et al. (2021)
Sunspot
S
Godahewa et al. (2021)
Appendix
15 Wu et al. Datasets
Wu & Keogh (2021)
Table 3: Anomaly detection baselines and datasets.
4.3 Forecasting
Baselines. In the main text, we compare against 12 baselines, with an additional 2 in Appendix A.4 (see Table 20). For the GPT2 specialist that we trained from scratch, we choose a lookback length of 96 for fair comparison with the other models in this paper. In total, we have 14 baselines. See Table 4A for a summary.
Datasets. For the in-domain testing regime, we test on 7 datasets, and for the zero-shot regime, we test on an additional 5. In total, we evaluate on 12 datasets. See Table 4B for a summary.
Metrics. We report the mean squared error (MSE) and mean absolute error (MAE) of the predicted versus the true forecast values.
A. Forecasting Baselines
Type
Model
Abbr.
Arch.
Citation
Multi-task
GPT2 - Generalist
GPT2
TF
Trained by us
GPT2 - Specialist
GPT2
TF
Trained by us w/ 96 lookback length
TimesNet
TiNet
Conv
Wu et al. (2022)
Single-task
iTransformer
iTrans
TF
Liu et al. (2023)
PatchTST
Patch
TF
Nie et al. (2022)
Crossformer
Cross
TF
Zhang & Yan (2022)
Fedformer
FED
TF
Zhou et al. (2022)
Non-stationary Trans.
Stat
TF
Liu et al. (2022b)
TiDE
TiDE
Das et al. (2023a)
RLinear
RLin
Linear
Li et al. (2023)
DLinear
DLin
Linear
Zeng et al. (2023)
SciNet
SCi
Liu et al. (2022a)
Appendix
N-Beats
Oreshkin et al. (2019)
MOMENT
Goswami et al. (2024)
B. Forecasting Datasets
Regime
Dataset
Abbr.
Citation
In-Domain
Weather
W
Liu et al. (2023)
Electricity
E
Liu et al. (2023)
Traffic
T
Liu et al. (2023)
ETTm1
m1
Liu et al. (2023)
ETTm2
m2
Liu et al. (2023)
ETTh1
h1
Liu et al. (2023)
ETTh2
h2
Liu et al. (2023)
Zero-Shot
Neuro2
N2
Peterson et al. (2022)
Neuro5
N5
Peterson et al. (2022)
Saugeen River Flow
R
Godahewa et al. (2021)
U.S. Births
B
Godahewa et al. (2021)
Sunspot
S
Godahewa et al. (2021)
Table 4: Forecasting baselines and datasets.
4.4 Task Selection
In the time series literature, there are five canonically studied tasks: imputation, anomaly detection, short- and long-term forecasting, and classification. In this work, we study imputation, anomaly detection, and long-term forecasting. We exclude short-term forecasting and classification for the following reasons.
Non-Standardized Baselines. The long-term forecasting task uses standardized input and output lengths across all datasets (in particular an input length of 96 timesteps and output lengths of 96, 192, 336, and 720 timesteps), as enforced by a large body of existing work Liu et al. (2023); Wu et al. (2022); Liu et al. (2022b); Zhou et al. (2022) among others [^2] Some methods utilize a $512$-dimensional input, which makes consistent comparisons challenging; despite this field-wide inconsistency, we include some of these results in Appendix 20. TOTEM outperforms other methods across lookback lengths $96,512$ at 58.3% AvgWins , while the next best model is GPT2 at 8.3% AvgWins .. This allows us to fairly baseline TOTEM without rerunning thousands of experiments on dozens of models trained from scratch.
In contrast, the short-term forecasting task typically uses non-standard and dataset-specific input and output dimensionalities (see Table 19 for details), which makes systematic, fair comparisons of TOTEM against prior works extremely challenging in the generalist setting [^3] Despite this, we show that TOTEM and GPT2 outperform all other methods on a small set of short term forecasting lengths and datasets in Appendix 18.. Thus, we exclude short-term forecasting from our main results.
Leaky Baselines. In both classification and anomaly detection, the modern SOTA baselines are leaky (Zhou et al., 2023; Wu et al., 2022; Xu et al., 2021), where leakage is defined as using the test set as the validation set during training. In particular, the cited works that report SOTA results all use models that were trained with either early stopping or with the best model checkpoint on the validation (i.e., the test) set. We felt strongly that we should not propagate faulty baselines, so we did not compare to these models in our work. Subsequent to the initial release of this paper, followup works have demonstrated on neural classification tasks that TOTEM, when compared to baselines trained in a non-leaky manner, achieves SOTA performance (Chau et al., 2024a; b).
For anomaly detection, the benchmark datasets used by Zhou et al. (2023); Wu et al. (2022); Xu et al. (2021) contain numerous flaws besides training leakage flawed (see Wu & Keogh (2021) for a detailed account). However, since Wu & Keogh (2021) released a large set of new, unflawed benchmarks, we elected to compare TOTEM to both the flawed and a subset of the unflawed baselines (see the comparisons to Wu & Keogh (2021) in the Appendix). Because we find that TOTEM convincingly achieves SOTA performance in both cases, we report our results to establish an unflawed baseline for future comparison.
In summary, due to non-standardized and leaky baselines, we only report systematic results on the imputation, anomaly detection, and long-term forecasting tasks.
5 Main Results
The primary goal of our experiments is to systematically evaluate TOTEM on multiple tasks simultaneously against new generalist benchmarks and strong specialist baselines (i.e., models trained on data from many domains versus one domain). In particular, for each task, we report evaluations against (i) specialists on the in-domain testing regime, (ii) generalists on the in-domain regime, and (iii) generalists on the zero-shot regime. We emphasize that no domain, sampling rate, or sensor dimension is shared between the training sets and zero-shot testing sets (see Table 6 for additional dataset details).
Throughout the main text, we report summary results. The full numerical results can be found throughout the Appendix. Moreover, all results are reported as the mean of 3 seeded runs, with standard deviations available in the Appendix. Since evaluation metrics differ across tasks, ($\downarrow$) will denote a metric where lower is better and ($\uparrow$) will denote a metric where higher is better. Given the varied metrics, we calculate the average number of best results, or AvgWins, for each method and highlight the best method. For a summary of training and testing domains, see Table 7; for a comparison of generalist parameter counts and training times, see Section A.8; for additional architecture and training details, see Sections A.11 and A.12.
[t] B. Generalist In-Domain Model TOTEM GPT2 Metric MSE MAE MSE MAE W 12.5% 0.029 0.060 0.029 0.045 25% 0.030 0.060 0.033 0.048 37.5% 0.032 0.062 0.037 0.054 50% 0.036 0.067 0.043 0.061 E 12.5% 0.065 0.171 0.080 0.186 25% 0.071 0.179 0.091 0.197 37.5% 0.080 0.189 0.108 0.213 50% 0.095 0.205 0.132 0.236 m1 12.5% 0.041 0.132 0.052 0.141 25% 0.044 0.135 0.065 0.154 37.5% 0.048 0.139 0.085 0.171 50% 0.058 0.152 0.117 0.196 m2 12.5% 0.040 0.125 0.029 0.095 25% 0.041 0.126 0.033 0.101 37.5% 0.043 0.129 0.038 0.110 50% 0.048 0.136 0.045 0.121 h1 12.5% 0.100 0.201 0.113 0.217 25% 0.108 0.209 0.131 0.231 37.5% 0.122 0.220 0.153 0.247 50% 0.144 0.237 0.182 0.266 h2 12.5% 0.075 0.175 0.067 0.155 25% 0.076 0.177 0.071 0.160 37.5% 0.093 0.195 0.077 0.167 50% 0.089 0.192 0.086 0.179 AvgWins 58.3% 43.8%
C. Generalist Zero-Shot
Model
TOTEM
GPT2
Metric
MSE
MAE
MSE
MAE
N2
12.5%
0.029
0.120
0.047
0.145
25%
0.033
0.127
0.064
0.164
37.5%
0.041
0.139
0.090
0.191
50%
0.056
0.160
0.131
0.228
N5
12.5%
0.017
0.085
0.021
0.095
25%
0.019
0.090
0.028
0.107
37.5%
0.022
0.098
0.039
0.123
50%
0.029
0.110
0.055
0.145
R
12.5%
0.071
0.109
0.093
0.119
25%
0.087
0.117
0.125
0.134
37.5%
0.112
0.129
0.167
0.154
50%
0.148
0.147
0.220
0.182
B
12.5%
0.632
0.642
0.392
0.496
25%
0.693
0.665
0.444
0.523
37.5%
0.761
0.692
0.498
0.553
50%
0.827
0.718
0.591
0.599
S
12.5%
0.057
0.160
0.070
0.173
25%
0.061
0.168
0.084
0.189
37.5%
0.069
0.178
0.103
0.209
50%
0.082
0.193
0.128
0.234
AvgWins
80.0% 20.0%
D. Specialist In-Domain